# On Resolution of $\ell_{1}$-norm minimization via a Two-metric Adaptive Projection Method

### v1

@@ 4.1 Global convergence
(18) also indicates the following:
$\mu_{k}\geq c\|x_{k}-\operatorname{prox}_{\gamma\|\cdot\|_{1}}(x_{k}-g_{k})\|^{\delta}.$
@@ 4.2 Manifold Identification and local superlinear convergence rate
$|x_{k}^{i}|=|x_{k}^{i}-\operatorname{prox}_{\gamma|\cdot|}(x_{k}^{i}-g_{k}^{i})|\leq\pi_{k}=\epsilon_{k}\quad\forall i\in\mathcal{A}(x^{*}),x_{k}\in B(x^{*},\delta_{1})$
Then,
$\mathcal{A}(x^{*})\subseteq I_{k}^{+}.$
There exists $\delta_{2}$, such that when $x_{k}\in B(x^{*},\delta_{2})$ and $\forall i\notin\mathcal{A}(x^{*})$,
@@
Then
$I_{k}^{+}\subseteq\mathcal{A}(x^{*}).$
Therefore, $I_{k}^{+}=\mathcal{A}(x^{*})$. Next we prove that $\mathcal{A}(x_{k+1})=I_{k}^{+}$. Indeed, let $b_{3}\leq\min\{\delta_{1},\delta_{2}\}$ such that for all $x_{k}\in B(x^{*},b_{3})$, $I_{k}^{+}=\mathcal{A}(x^{*})$ and $\|\bar{g}_{k}+\bar{\omega}_{k,\epsilon}\|$ small enough. Combine Eq. 15 with the fact that $\|\bar{g}_{k}+\bar{\omega}_{k,\epsilon}\|$ small enough and $t_{k}\leq 1$, we have $t_{k}\|\bar{p}_{k}\|$ can be arbitary small.
@@
we have that $x_{k+1}^{i}=0$ by the update rule. Note that when $b_{3}$ is small enough, right-hand side in Eq. 20 is less than any constant. Therefore, we only need to argue that for any $i\in I_{k}^{+}$
@@
$I_{k}^{+}$ =$=\left\{i\mid|x_{k}^{i}|\leq\epsilon_{k},|g_{k}^{i}|<\gamma\right\}$
$I_{k}^{-+}$ =$=\left\{i\mid x_{k}^{i}>\epsilon_{k}\right\}$
$I_{k}^{--}$ =$=\left\{i\mid x_{k}^{i}<-\epsilon_{k}\right\}.$
There exists a neighborhood $B(x^{*},\bar{b}_{3})\subseteq B(x^{*},b_{3})$ such that
$\|\bar{p}_{k}\|\leq\frac{b_{3}}{2}.$
Therefore, for all $x_{k}\in B(x^{*},\min\{\frac{b_{3}}{2},\bar{b}_{3}\})$,
@@
Let $\tilde{X}^{*}:=X^{*}\cap\mathcal{M}_{*}$ and define $\Pi_{\tilde{X}^{*}}(x)=\operatorname{argmin}_{y\in\tilde{X}^{*}}\|y-x\|$. Pick any $\bar{x}_{k}\in\Pi_{\tilde{X}^{*}}(x_{k})$ and note that $\left\|\bar{x}_{k}-x^{*}\right\|\leq\left\|\bar{x}_{k}-x_{k}\right\|+\left\|x_{k}-x^{*}\right\|\leq 2\left\|x_{k}-x^{*}\right\|\leq b_{2}$. Therefore, we can apply 2.12 and results in Lemma 2.4. Note that ${\rm grad}_{\mathcal{M}_{*}}\psi(\bar{x}_{k})=0$ and $[\operatorname{Hess}_{\mathcal{M}_{*}}\psi(x_{k})(x_{k}-\bar{x}_{k})]_{I_{k}^{-}}=H_{k}(x_{k}-\bar{x}_{k})_{I_{k}^{-}}$ since $x_{k}$ and $\bar{x}_{k}$ are both on $\mathcal{M}_{*}$. Therefore,
@@
$\|(H_{k}+\mu_{k}I)^{-1}H_{k}\|=\|I-(H_{k}+\mu_{k}I)^{-1}\mu_{k}I\|\leq\left(\|I\|+\left\|\left(H_{k}+\mu_{k}I\right)^{-1}\mu_{k}I\right\|\right)\leq 2$
@@
$\leq\left(\frac{1}{2}L_{H}c_{1}^{\delta}c^{-1}+2\right)\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right)$
@@
$\leq$ $\frac{L_{H}}{6}\left\|s_{k}\right\|^{3}-\langle s_{k},\operatorname{grad}_{\mathcal{M}_{*}}f(x_{k})\rangle+\frac{1}{2}\langle s_{k},\operatorname{Hess}_{\mathcal{M}_{*}}f(x_{k})[s_{k}]\rangle+\bar{\omega}_{k,\epsilon}^{T}\left(\hat{x}^{k+1}-x^{k}\right)$
@@
$\frac{L_{H}}{6}\left\|\bar{p}_{k}\right\|-(1-\tau)\mu_{k}\leq-\sigma(1-\tau)\mu_{k}$
i.e.
$\frac{L_{H}}{6}\left\|\bar{p}_{k}\right\|\leq(1-\sigma)(1-\tau)\mu_{k}$
Note that from Lemma 4.14 and Eq. 21, for $\delta\in(0,1)$, this sufficient condition can be satisfied when $b_{5}$ is small enough since $\|\bar{g}_{k}+\bar{\omega}_{k,\epsilon}\|$ can be arbitary small.
@@
Note that $x_{k+1}=\hat{x}_{k+1}$ defined in Eq. 22, and
$\|\operatorname{grad}_{\mathcal{M}_{*}}\psi(x_{k})-\operatorname{grad}_{\mathcal{M}_{*}}\psi(x_{k+1})-\operatorname{Hess}_{\mathcal{M}_{*}}\psi(x_{k})(x_{k}-x_{k+1})\|\leq\frac{1}{2}L_{H}\|\bar{p}_{k}\|^{2}$
and
$\mu_{k}=c\|\operatorname{grad}_{\mathcal{M}_{*}}\psi(x_{k})\|^{\delta}=c\|\operatorname{grad}_{\mathcal{M}_{*}}\psi(x_{k})-\operatorname{grad}_{\mathcal{M}_{*}}\psi(\bar{x}_{k})\|^{\delta}\leq cL_{\psi}^{\delta}\|x_{k}-\bar{x}_{k}\|^{\delta}$
Therefore, combine this with EB and Lemma 4.14, we have
@@
$=\mathcal{O}(\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right))^{1+\delta}+\mathcal{O}(\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right))^{2}$
$=\mathcal{O}(\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right))^{1+\delta}$
@@
$r=\min\{\frac{1}{2}c_{3}^{-\frac{1}{\delta}},\frac{b_{6}}{1+c_{2}+2c_{2}\frac{2^{-\delta}}{1-2^{-\delta}}}\}<b_{6},$
@@
$\|x_{k+1}-x^{*}\|\leq\|x_{k}-x^{*}\|+\|\bar{p}_{k}\|\leq r+c_{2}\|x_{k}-\bar{x}_{k}\|\leq r+c_{2}r\leq b_{6}$
Suppose that $x_{k+j}\in B(x^{*},b_{6})$ for $1\leq j\leq i$, then from Lemma 4.18, let $q=1+\delta$ we have
$\|x_{k+j}-\bar{x}_{k+j}\|\leq c_{3}\|x_{k+j-1}-\bar{x}_{k+j-1}\|^{q}\leq\ldots\leq c_{3}^{\frac{1-q^{j}}{1-q}}\|x_{k}-\bar{x}_{k}\|^{q^{j}}\leq c_{3}^{\frac{1-q^{j}}{1-q}}r^{q^{j}}\leq 2r2^{-q^{j}}$
Note that
@@
$\leq r+c_{2}r+2c_{2}r\sum_{j=1}^{i}2^{-q^{j}}$
Since
$(1+\delta)^{j}\geq 1+\delta j\geq\delta j$
Therefore
$\sum_{j=1}^{i}2^{-q^{j}}\leq\sum_{j=1}^{\infty}2^{-\delta j}$
And
$\|x_{k+i+1}-x^{*}\|\leq r+c_{2}r+2c_{2}r\sum_{j=1}^{\infty}2^{-\delta j}\leq r+c_{2}r+2c_{2}r\frac{2^{-\delta}}{1-2^{-\delta}}\leq b_{6}$
@@
From manifold EB and Theorem 4.7, we have
$\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right)\rightarrow 0.$
Thus, for $\epsilon>0$ and $\rho<1$, there exists $K$ such that for all $k\geq K$, we have
$\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right)<\epsilon\quad\text{and}\quad\operatorname{dist}\left(x_{k+1},X^{*}\cap\mathcal{M}_{*}\right)\leq\rho\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right)$
@@
Together with Lemma 4.18, we have
@@ 5 Numerical experiment
In this section, we study the practical performance of the two-metric adaptive projection method and compare it with some existing algorithms. All experiments are coded in MATLAB (R2024a) and run on a desktop with a 13600KF CPU and 32 GB of RAM. We will compare the performance of the two-metric adaptive projection method with some state-of-the-art algorithms on logistic regression problems and LASSO problems.
@@ 5.1 Numerical results on logistic regression problem
Note that A is sparse in many real data sets, therefore $\nabla^{2}f(x)p$ can be computed efficiently with worst case complexity $\mathcal{O}(nm)$.
@@
Alternating Newton acceleration
This is an accelerated proximal gradient algorithm [3], which alternates between proximal gradient step and Riemannian Newton step on the current manifold. It has the ability for manifold identification.
IRPN
This is the family of inexact SQA methods proposed in [30], and we use the same setting of this paper, i.e. $\theta=\beta=0.25,\zeta=0.4,c=10^{-6}$ and $\eta=0.5$. We choose $\rho=0.5$ since IRPN has superlinear convergence in this setting and performs best in the paper. The coordinate descent algorithm is used for inner problem of IRPN.
Two metric adaptive projection
This is Algorithm 1 we proposed in this paper. We set $c=10^{-4}$, $\tau=10^{-1}$ and $\delta=\frac{1}{2}$.
@@
While SQA type methods inherently possess manifold identification capabilities, they fail to leverage this property in their algorithmic design, resulting in higher computational costs. In contrast, although AltN is suboptimal for manifold identification, it demonstrates faster performance than IRPN for some datasets. To address this gap, ISQA+, a hybrid method proposed in [15], integrates the strengths of SQA and alternating Newton. Specifically, it employs SQA for initial manifold identification and subsequently switches to an alternating Newton method for accelerated optimization.
@@ 5.2 Numerical results on LASSO problem
Here, $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$ are given data samples; $\gamma$ is a given regularization parameter. Closely related to the LASSO problem is the Basis Pursuit problem, which is a convex optimization problem that seeks the sparsest solution to an underdetermined linear system of equations. The Basis Pursuit (BP) problem can be formulated as
@@
The theory for penalty functions implies that the solution of the LASSO problem goes to the solution of BP problem as $\gamma$ goes to zero.
In our experiment, we compare our algorithm with other state-of-the-art algorithms on large-scale reconstruction problems. The test problems are from [28, 17], which are generated by the following procedure. Firstly, we randomly generate a sparse solution $\bar{x}\in\mathbb{R}^{n}$ with $k$ nonzero entries, where $n=512^{2}=262144$ and $k=\left\lfloor n/40\right\rfloor=6553$. The $k$ different indices are uniformly chosen from $\{1,2,\ldots,n\}$ and we set the magnitude of each nonzero element by $\bar{x}_{i}=\eta_{1}(i)10^{d\eta_{2}(i)/20}$, where $\eta_{1}(i)$ is randomly chosen from $\{-1,1\}$ with probability $\frac{1}{2}$, respectively, $\eta_{2}(i)$ is uniformly distributed in $[0,1]$ and $d$ is a dynamic range which can influence the efficiency of the solvers. Then we choose $m=n/8=32768$ random cosine measurements, i.e., $Ax=(dct(x))_{J}$, where $J$ contains $m$ different indices randomly chosen from $\{1,2,\ldots,n\}$ and $dct$ is the discrete cosine transform. Finally, the input data is specified by $b=A\bar{x}+\epsilon$, where components of $\epsilon$ are i.i.d Gaussian noises with a standard deviation $\bar{\sigma}=0.1$.

### v2

@@ 4.1 Global convergence
(17) also indicates the following: $\mu_{k}\geq c\|x_{k}-\operatorname{prox}_{\gamma\|\cdot\|_{1}}(x_{k}-g_{k})\|^{\delta}.$
@@ 4.2 Manifold Identification and local superlinear convergence rate
$|x_{k}^{i}|=|x_{k}^{i}-\operatorname{prox}_{\gamma|\cdot|}(x_{k}^{i}-g_{k}^{i})|\leq\pi_{k}=\epsilon_{k},\quad\forall i\in\mathcal{A}(x^{*}),x_{k}\in B(x^{*},\delta_{1}).$
Then $\mathcal{A}(x^{*})\subseteq I_{k}^{+}.$ There exists $\delta_{2}$, such that when $x_{k}\in B(x^{*},\delta_{2})$ and $\forall i\notin\mathcal{A}(x^{*})$,
@@
Then $I_{k}^{+}\subseteq\mathcal{A}(x^{*}).$ Therefore, $I_{k}^{+}=\mathcal{A}(x^{*})$. Next we prove that $\mathcal{A}(x_{k+1})=I_{k}^{+}$. Indeed, let $b_{3}\leq\min\{\delta_{1},\delta_{2}\}$ such that for all $x_{k}\in B(x^{*},b_{3})$, $I_{k}^{+}=\mathcal{A}(x^{*})$ and $\|\bar{g}_{k}+\bar{\omega}_{k,\epsilon}\|$ small enough. Combine Eq. 14 with the fact that $\|\bar{g}_{k}+\bar{\omega}_{k,\epsilon}\|$ small enough and $t_{k}\leq 1$, we have $t_{k}\|\bar{p}_{k}\|$ can be arbitary small.
@@
we have that $x_{k+1}^{i}=0$ by the update rule. Note that when $b_{3}$ is small enough, right-hand side in Eq. 19 is less than any constant. Therefore, we only need to argue that for any $i\in I_{k}^{+}$,
@@
$I_{k}^{+}=\left\{i\mid|x_{k}^{i}|\leq\epsilon_{k},|g_{k}^{i}|<\gamma\right\},\;I_{k}^{-+}=\left\{i\mid x_{k}^{i}>\epsilon_{k}\right\},\;I_{k}^{--}=\left\{i\mid x_{k}^{i}<-\epsilon_{k}\right\}.$
There exists a neighborhood $B(x^{*},\bar{b}_{3})\subseteq B(x^{*},b_{3})$ such that $\|\bar{p}_{k}\|\leq\frac{b_{3}}{2}.$ Therefore, for all $x_{k}\in B(x^{*},\min\{\frac{b_{3}}{2},\bar{b}_{3}\})$,
@@
Let $\tilde{X}^{*}:=X^{*}\cap\mathcal{M}_{*}$ and define
(21) $\Pi_{\tilde{X}^{*}}(x)\triangleq\operatorname{argmin}_{y\in\tilde{X}^{*}}\|y-x\|,\;\bar{x}_{k}=\Pi_{\tilde{X}^{*}}(x_{k}).$
Note that $\left\|\bar{x}_{k}-x^{*}\right\|\leq\left\|\bar{x}_{k}-x_{k}\right\|+\left\|x_{k}-x^{*}\right\|\leq 2\left\|x_{k}-x^{*}\right\|\leq b_{2}$. Therefore, we can apply 2.12 and results in Lemma 2.4. Note that ${\rm grad}_{\mathcal{M}_{*}}\psi(\bar{x}_{k})=0$ and $[\operatorname{Hess}_{\mathcal{M}_{*}}\psi(x_{k})(x_{k}-\bar{x}_{k})]_{I_{k}^{-}}=H_{k}(x_{k}-\bar{x}_{k})_{I_{k}^{-}}$ since $x_{k}$ and $\bar{x}_{k}$ are both on $\mathcal{M}_{*}$. Therefore,
@@
$\|(H_{k}+\mu_{k}I)^{-1}H_{k}\|=\|I-(H_{k}+\mu_{k}I)^{-1}\mu_{k}I\|\leq\left(\|I\|+\left\|\left(H_{k}+\mu_{k}I\right)^{-1}\mu_{k}I\right\|\right)\leq 2.$
@@
$\leq\left(\frac{1}{2}L_{H}c_{1}^{\delta}c^{-1}+2\right)\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right).$
@@
$\leq$ $\frac{L_{H}}{6}\left\|s_{k}\right\|^{3}-\langle s_{k},\operatorname{grad}_{\mathcal{M}_{*}}f(x_{k})\rangle+\frac{1}{2}\langle s_{k},\operatorname{Hess}_{\mathcal{M}_{*}}f(x_{k})[s_{k}]\rangle+\omega_{k,\epsilon}^{T}\left(\hat{x}^{k+1}-x^{k}\right)$
@@
$\frac{L_{H}}{6}\left\|\bar{p}_{k}\right\|-(1-\tau)\mu_{k}\leq-\sigma(1-\tau)\mu_{k}\Longleftrightarrow\frac{L_{H}}{6}\left\|\bar{p}_{k}\right\|\leq(1-\sigma)(1-\tau)\mu_{k}.$
Note that from Lemma 4.14 and Eq. 20, for $\delta\in(0,1)$, this sufficient condition can be satisfied when $b_{5}$ is small enough.
@@
Note that we have $x_{k+1}=\hat{x}_{k+1}$ defined in Eq. 22. Moreover,
$\|\operatorname{grad}_{\mathcal{M}_{*}}\psi(x_{k})-\operatorname{grad}_{\mathcal{M}_{*}}\psi(x_{k+1})-\operatorname{Hess}_{\mathcal{M}_{*}}\psi(x_{k})(x_{k}-x_{k+1})\|\leq\frac{1}{2}L_{H}\|\bar{p}_{k}\|^{2},$
$\mu_{k}=c\|\operatorname{grad}_{\mathcal{M}_{*}}\psi(x_{k})\|^{\delta}=c\|\operatorname{grad}_{\mathcal{M}_{*}}\psi(x_{k})-\operatorname{grad}_{\mathcal{M}_{*}}\psi(\bar{x}_{k})\|^{\delta}\leq cL_{\psi}^{\delta}\|x_{k}-\bar{x}_{k}\|^{\delta},$
where $\bar{x}_{k}$ is defined as in (21). Therefore, combine the above inequalities with manifold EB, Lemma 4.14, and definition of $r_{k}$, we have
@@
$\leq\mathcal{O}(\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right))^{1+\delta}+\mathcal{O}(\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right))^{2}$
$\leq\mathcal{O}(\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right))^{1+\delta}.$
@@
$r\triangleq\min\left\{\frac{1}{2}c_{3}^{-\frac{1}{\delta}},\frac{b_{6}}{1+c_{2}+2c_{2}\frac{2^{-\delta}}{1-2^{-\delta}}}\right\}<b_{6},$
@@
$\|x_{k+1}-x^{*}\|\leq\|x_{k}-x^{*}\|+\|\bar{p}_{k}\|\leq r+c_{2}\|x_{k}-\bar{x}_{k}\|\leq r+c_{2}r\leq b_{6}.$
Suppose $x_{k+j}\in B(x^{*},b_{6})$ for $1\leq j\leq i$. From Lemma 4.18 (let $q=1+\delta$) we have
$\|x_{k+j}-\bar{x}_{k+j}\|\leq c_{3}\|x_{k+j-1}-\bar{x}_{k+j-1}\|^{q}\leq\ldots\leq c_{3}^{\frac{1-q^{j}}{1-q}}\|x_{k}-\bar{x}_{k}\|^{q^{j}}\leq c_{3}^{\frac{1-q^{j}}{1-q}}r^{q^{j}}\leq 2r2^{-q^{j}},$
where $\bar{x}_{k}$ is defined as in (21) and the last inequality is by the definition of $r$. Note
@@
(23) $\leq r+c_{2}r+2c_{2}r\sum_{j=1}^{i}2^{-q^{j}}.$
By noticing
$(1+\delta)^{j}\geq 1+\delta j\geq\delta j\implies\sum_{j=1}^{i}2^{-q^{j}}\leq\sum_{j=1}^{\infty}2^{-\delta j},$
and combine (23), we have
$\|x_{k+i+1}-x^{*}\|\leq r+c_{2}r+2c_{2}r\sum_{j=1}^{\infty}2^{-\delta j}\leq r+c_{2}r+2c_{2}r\frac{2^{-\delta}}{1-2^{-\delta}}\leq b_{6},$
@@
From manifold EB and Theorem 4.7, we have $\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right)\rightarrow 0.$ Thus, for $\epsilon>0$ and $\rho<1$, there exists $K$ such that for all $k\geq K$, we have $x_{k}\in B(x^{*},b_{6})$, and
$\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right)<\epsilon\quad\text{and}\quad\operatorname{dist}\left(x_{k+1},X^{*}\cap\mathcal{M}_{*}\right)\leq\rho\operatorname{dist}\left(x_{k},X^{*}\cap\mathcal{M}_{*}\right).$
@@
Together with Lemma 4.18, we have for all $k\geq K$,
@@ 5 Numerical experiment
In this section, we study the practical performance of the two-metric adaptive projection method and compare it with some existing algorithms. All experiments are coded in MATLAB (R2024a) and run on a desktop with a 13600KF CPU and 32 GB of RAM. The source code can be downloaded from https://github.com/Hanju-Wu/TMAP. We will compare the performance of the two-metric adaptive projection method with some state-of-the-art algorithms on logistic regression problems and LASSO problems.
@@ 5.1 Numerical results on logistic regression problem
Note that A is sparse in many real data sets, therefore $\nabla^{2}f(x)p$ can be computed efficiently with worst-case complexity $\mathcal{O}(nm)$.
@@
  1. 1.
    Alternating Newton acceleration (AltN) This is an accelerated proximal gradient algorithm [3], which alternates between proximal gradient step and Riemannian Newton step on the current manifold. It has the ability for manifold identification.
  2. 2.
    IRPN This is the family of inexact SQA methods proposed in [35], and we use the same setting of this paper, i.e. $\theta=\beta=0.25,\zeta=0.4,c=10^{-6}$ and $\eta=0.5$. We choose $\rho=0.5$ since IRPN has superlinear convergence in this setting and performs best in the paper. The coordinate descent algorithm is used for inner problem of IRPN.
  3. 3.
    Two metric adaptive projection (TMAP) This is Algorithm 1 we proposed in this paper. We set $c=10^{-4}$, $\tau=10^{-1}$ and $\delta=\frac{1}{2}$.
@@
While SQA type methods inherently possess manifold identification capabilities, they fail to leverage this property in their algorithmic design, resulting in higher computational costs. In contrast, although AltN is suboptimal for manifold identification, it shows better performance than IRPN for some datasets. To address this gap, ISQA+, a hybrid method proposed in [16], integrates the strengths of SQA and alternating Newton. Specifically, it employs SQA for initial manifold identification and subsequently switches to an alternating Newton method for accelerated optimization.
@@
Also note that the convergence behavior of SQA type methods in numerical experiments is not linearly dependent on the maximum number of inner iterations and accuracy of inner iterations. Fast convergence requires careful tuning of the inner solver accuracy in practice.
@@ 5.2 Numerical results on LASSO problem
Here, $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$ are given data samples; $\gamma$ is a given regularization parameter. Closely related to the LASSO problem is the Basis Pursuit problem which is a convex optimization problem that seeks the sparsest solution to an underdetermined linear system of equations. The Basis Pursuit (BP) problem can be formulated as
@@
The theory for penalty functions implies that the solution of the LASSO problem goes to the solution of BP problem as $\gamma$ goes to zero. Numerical experiments suggest that in practice, most S-sparse signals are in fact recovered exactly via BP once $m\geq 4S$ or so and using Fourier measurements [10].
In our experiment, we compare our algorithm with other state-of-the-art algorithms on large-scale reconstruction problems (LASSO with $\gamma\approx 0.01$). The test problems are from [33, 21], which are generated by the following procedure. Firstly, we randomly generate a sparse solution $\bar{x}\in\mathbb{R}^{n}$ with $k$ nonzero entries, where $n=512^{2}=262144$ and $k=\left\lfloor n/40\right\rfloor=6553$. The $k$ different indices are uniformly chosen from $\{1,2,\ldots,n\}$ and we set the magnitude of each nonzero element by $\bar{x}_{i}=\eta_{1}(i)10^{d\eta_{2}(i)/20}$, where $\eta_{1}(i)$ is randomly chosen from $\{-1,1\}$ with probability $\frac{1}{2}$, respectively, $\eta_{2}(i)$ is uniformly distributed in $[0,1]$ and $d$ is a dynamic range which can influence the efficiency of the solvers. Then we choose $m=n/8=32768$ random cosine measurements, i.e., $Ax=(dct(x))_{J}$, where $J$ contains $m$ different indices randomly chosen from $\{1,2,\ldots,n\}$ and $dct$ is the discrete cosine transform. Finally, the input data is specified by $b=A\bar{x}+\epsilon$, where components of $\epsilon$ are i.i.d Gaussian noises with a standard deviation $\bar{\sigma}=0.1$.
@@
Another efficient solver for LASSO is the SSNAL [20]. The authors demonstrate that SSNAL achieves superior computational efficiency for large-scale regression problems with real-world datasets. This advantage stems from the fact that the real-world datasets they use usually has moderate sample size far less than the parameters ($m\ll n$). SSNAL solves the dual problem of LASSO instead of the primal problem, which circumvents numerical instabilities caused by ill-conditioned primal problems when matrix $A$ exhibits poor spectral properties. Furthermore, solving Newton linear system from the dual formulation costs $\mathcal{O}(\min\{m^{2}(m+r),r^{2}(m+r)\})$. Note that we usually have $r\ll n$, especially when the regularizer coefficient $\gamma$ is chosen aggressively.
In contrast, for synthetic reconstruction problems (e.g., compressed sensing with random Gaussian matrices or random discrete cosine transform), the sensing matrix $A$ typically has a significantly higher number of rows than the matrices in regression tasks in [20]. Also, the choice of $\gamma$ is usually small compared to the regression tasks. We compare our algorithm with SSNAL on the reconstruction problem with random Gaussian matrices and discrete cosine transform (DCT) matrices with sparse signals (sparsity level $m/5$). The results are presented in Table 6.
Table 6: Computational results on LASSO problem with SSNAL and TMAP. The CPU time (in seconds) is averaged over 10 independent runs. NaN means that SSNAL fails to solve the problem within a reasonable time limit.
random matrix $A$ Size TMAP SSNAL
DCT $n=100^{2},m=n/8,\gamma=0.01$ 0.01 2.21
$n=200^{2},m=n/8,\gamma=0.01$ 0.06 27.46
Gaussian $n=1000,m=500,\gamma=0.001$ 0.14 0.46
$n=2000,m=1000,\gamma=0.001$ 0.63 1.35
$n=4000,m=2000,\gamma=0.001$ 2.90 4.62
$n=8000,m=4000,\gamma=0.001$ 15.26 NaN